What is the concentration of the H+ ions in a 0.001N KOH solution

a)10^-17
b)10^-11
c)10^-10
d)10^-3
1
Expert's answer
2017-10-20T05:00:06-0400
In solution:
K_w=[H^+ ]×[〖OH〗^- ]=10^(-14)
[H^+ ]=(10^(-14))/[〖OH〗^- ]
For the KOH, 0.001 N = 0.001 M
KOH (aq) = K+ (aq) + OH- (aq)
[〖OH〗^- ]=c(KOH)=0.001 M
[H^+ ]=(10^(-14))/(0.001 M)=10^(-11) M
Answer is b) 10-11